OpenStudy (anonymous):

Apollo 11 was launched at 9:30 a.m. eastern daylight time on july 16,1969. The mission lasted 195h 18min. what was the date and time when the mission ended?

OpenStudy (anonymous):

need help!!!!!!!!

OpenStudy (gw2011):

I got an answer of: July 24, 1969 at 12:48 p.m. Here is how I arrived at it: 195h/24h=8 days (8days)(24 hours)=192 hours 195-192=3 hours remaining plus the 18 min July 16+8 days=July 24 9:30 a.m. + 3 hours and 18 min left = 12:48 p.m. Therefore, it is July 24, 1969 at 12:48 p.m.